a group of 5 cards is numbered 1-5 you choose one card at random. find each theoretical probability?
part a
P(even number)
part b
P(Prime Number)
part c
P(less than 5)
part d
P(card is 2)
thanks.

Answer by fcabanski(1391)

Probability is desired/total. There are 5 total cards, so that's the denominator.


Even number - there are 2 even numbers so P(even) = 2/5 = 40%.
Prime number - there are 3 prime numbers (2, 3, 5) so P(prime) = 3/5 = 60%.
Less than 5 - there are 4 less than 5 (1, 2, 3, 4) so P(less than 5) = 4/5=80%.
2 - there is one 2 so P(2) = 1/5=20%.

If you need help understanding math so you can solve these problems
yourself, then one on one online tutoring is the answer ($30/hr). If you
need faster solutions with guaranteed detailed answers, then go with
personal problem solving ($3.50-$5.50 per problem). Contact me at
fcabanski@hotmail.com